OpenStudy (anonymous):

need help with 3 through 20 Mostly explaining how to do the work. glencoe.com/sites/common_assets/mathematics/alg2_2010/worksheets/alg2hwp_72.pdf

jimthompson5910 (jim_thompson5910):

In #3, are there any portions where you're dividing by a variable?

OpenStudy (anonymous):

what?

jimthompson5910 (jim_thompson5910):

look at #3 do you see any divisions by a variable

jimthompson5910 (jim_thompson5910):

ex: 2/x is the number 2 divided by a variable

OpenStudy (anonymous):

yes 2/m^2

jimthompson5910 (jim_thompson5910):

so that invalidates it from being a polynomial

jimthompson5910 (jim_thompson5910):

ie 2/m^2 is NOT a polynomial

jimthompson5910 (jim_thompson5910):

because you can't divide by a variable if you want a polynomial

OpenStudy (anonymous):

oh ok thx

jimthompson5910 (jim_thompson5910):

for number 9, you replace every x with -2 and then you evaluate \[\Large p(x) = x^4 + \frac{1}{2}x^3 - \frac{1}{2}x\] \[\Large p(-2) = (-2)^4 + \frac{1}{2}(-2)^3 - \frac{1}{2}(-2)\] \[\Large p(-2) = 16 + \frac{1}{2}(-8) - \frac{1}{2}(-2)\] \[\Large p(-2) = 16 - 4 + 1\] \[\Large p(-2) = 13\]

OpenStudy (anonymous):

ok thx